OpenStudy (nerdy_3000):

How do I solve this using the(i=-1) or i=square root -1 -2(x+1)^2=72

OpenStudy (anonymous):

divide by\(-2\) then take the square root

OpenStudy (anonymous):

\[\frac{72}{-2}=-36\] and \[\sqrt{-36}=\sqrt{36\times (-1)}=\sqrt{36}\sqrt{-1}=6i\]

OpenStudy (nerdy_3000):

Thank you Satellite73!

OpenStudy (anonymous):

yw don't forget to subtract \(1\) to get your final answer